OpenStudy (anonymous):

6log(c )z-1/2log(c ) x+log(c )w

jimthompson5910 (jim_thompson5910):

each log has a base of c?

OpenStudy (anonymous):

i will draw it

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

you can always use latex or the equation editor =] @c_lapham

OpenStudy (anonymous):



OpenStudy (anonymous):

says to write as single logarithm

OpenStudy (anonymous):

Is it:\[\bf 6\log _{c}(z)-\frac{ 1 }{ 2 }\log_c(x)+\log_c(w)\]@c_lapham

OpenStudy (anonymous):

@c_lapham

OpenStudy (anonymous):

yes but there isn't parenthesis on the variables if that makes a difference

OpenStudy (anonymous):

Notice that the coefficients in front of the logarithms can be made exponents. So we can re-write the expression like this:\[\bf 6\log_c(z)-\frac{1}{2}\log_c(x)+\log_c(w)=\log_c(z)^6-\log_c(x)^{1/2}+\log_c(w)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now use the following rules to write as a single expression.\[\bf \log_c(a \times b)=\log_c(a)+\log(b)\]\[\bf \log_c(a \div b)=\log_c(a)-\log_c(b)\]Then you'll get the answer. @c_lapham I have to leave bye.

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

\[ \log_{c}(z)^{6}\div (x)^{1/2 }\]

OpenStudy (anonymous):

would this be right for the first half?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

what about the last log set what would i do with that

jimthompson5910 (jim_thompson5910):

well you now have something like this log(....) + log(....) or something like this log(A) + log(B) so you would combine the two logs to get log(A*B)

OpenStudy (anonymous):

but isnt there only one log left to simplify

jimthompson5910 (jim_thompson5910):

doing that will give you \[\large \log_{c}\left(\frac{z^6}{x^{1/2}}*w\right)\] which can be rewritten to get \[\large \log_{c}\left(\frac{wz^6}{x^{1/2}}\right)\]

OpenStudy (anonymous):

oh i see what you did

jimthompson5910 (jim_thompson5910):

\[\large \log_{c}\left(\frac{z^6}{x^{1/2}}\right) + \log_{c}\left(w\right)\] would turn into \[\large \log_{c}\left(\frac{z^6}{x^{1/2}}*w\right)\] after using the rule: log(A)+log(B) = log(A*B)

OpenStudy (anonymous):

thank you i actually understand it now

jimthompson5910 (jim_thompson5910):

you're welcome